# Diffuse-CLoC: Guided Diffusion for Physics-based Character Look-ahead Control

### v2

@@ V-F Effect of Different Rolling Schemes
Shown in Table II, we observe a trade-off in rolling schemes across tasks. In the jump task, disabling state rolling causes changes in the motion plan which disrupt the character’s momentum, leading to lower jump heights and poor success rates. Conversely, in the forest task, an optimal plan for the current timestep may become suboptimal as distant obstacles enter the planning horizon. Enforcing consistency with these suboptimal plans lowers success rates and extends completion time. Despite these trade-offs, our best model performs robustly across scenarios with a single set of parameters. For action rolling, the impact on performance is minimal, enabling a $25\%$ acceleration in the diffusion process without significant loss.
VI Discussion and Future Work
@@
In this work, we introduced Diffuse-CLoC, a guided diffusion model for physics-based character look-ahead control. By co-diffusing a long horizon of states and actions, our approach overcomes limitations in existing methods, such as limited steerability and physical inconsistencies in generated motions. Diffuse-CLoC employs inference-time conditioning techniques to leverage intuitive state-space guidance for generation in action space, enabling flexible reuse for novel downstream tasks without retraining or finetuning. Experimental results highlight its zero-shot capability to address diverse tasks, including static and dynamic obstacle avoidance, dynamic maneuvers, and task-space control using a single model.
@@
Our method has some limitations. The guidance strength needs careful tuning to balance task completion while staying close to the training distribution. Future work could explore alternative guidance formulation, such as using future returns instead of immediate rewards or trust region constraints [43]. Additionally, our method’s reliance on historical observations biases it toward seen motion patterns, limiting transitions as noted in [26]. For example, some characters in the video are not very responsive to the guided signals in their tasks. We believe historical observations impose a constraint on the character, making the effect of future state predictions propagate slower to affect the next action. Future work could explore reducing this dependency. Futhermore, some results show lower motion quality, like foot jitter, likely caused by torque noise added during data augmentation. Without this noise, diffusion fails due to compounding errors [7]. Future work could explore alternative augmentation methods to reduce compounding errors while improving motion smoothness.

### v3

@@ 5.6. Effect of Different Rolling Schemes
Shown in Table 2, we observe a trade-off in rolling schemes across tasks. In the jump task, disabling state rolling causes changes in the motion plan which disrupt the character’s momentum, leading to lower jump heights and poor success rates. Conversely, in the forest task, an optimal plan for the current timestep may become suboptimal as distant obstacles enter the planning horizon. Enforcing consistency with these suboptimal plans lowers success rates and extends completion time. Despite these trade-offs, our best model performs robustly across scenarios with a single set of parameters. For action rolling, the impact on performance is minimal, enabling a $25\%$ acceleration in the diffusion process without significant loss.
Figure 8. A long-horizon sequence of diverse tasks. A sequence of tasks in a single run, including route path following, platform jumping, and reaching a waypoint behind cylindrical barriers.
6. Discussion and Future Work
@@
In this work, we introduced Diffuse-CLoC, a guided diffusion model for physics-based character look-ahead control. By co-diffusing a long horizon of states and actions, our approach overcomes limitations in existing methods, such as limited steerability and physical inconsistencies in generated motions. Diffuse-CLoC employs inference-time conditioning techniques to leverage intuitive state-space guidance for generation in action space, enabling flexible reuse for novel downstream tasks without retraining or finetuning. Experimental results highlight its zero-shot capability to address diverse tasks, including static and dynamic obstacle avoidance, dynamic maneuvers, and task-space control using a single model. We showcase a long-horizon run with these diverse tasks in Fig. 8.
Despite its effectiveness, our approach has several limitations that open directions for future research. One key challenge is tuning the strength of classifier guidance to ensure task success while staying within the training distribution. For some tasks, particularly those involving jumping or crawling, overly strong guidance sometimes led to unnatural motions or degraded quality. In these cases, increased guidance weight on obstacle clearance was necessary to ensure successful execution. Exploring alternative guidance formulations such as using future returns instead of immediate rewards or incorporating trust region constraints (Huang et al., 2024b) could provide better control.
Another limitation stems from insufficient data coverage. In tasks like hand-reaching or root path tracking, the model sometimes exhibits erroneous behaviors. For instance, jumping in place when asked to reach overhead. This is because jumping motions are present in the training dataset while overhead-reaching motions are not. While additional data would help mitigate these issues, our experiments were designed to evaluate performance under limited coverage.
We also observed reduced motion quality in some results, such as foot jitter. This is likely caused by torque noise added during data augmentation. However, without this noise, diffusion fails due to compounding errors (Truong et al., 2024). Future work could explore alternative augmentation methods to reduce compounding errors while improving motion smoothness.
@@
In addition, the model’s reliance on historical observations biases it toward previously seen motion patterns, which can limit transitions and responsiveness, as noted in (Chen et al., 2024). For example, some characters in the video are not very responsive to the guided signals in their tasks. We believe historical observations impose a constraint on the character, making the effect of future state predictions propagate slower to affect the next action. Future work could explore reducing this dependency.
Lastly, extending our approach to human-object interaction remains an open challenge. Physical manipulation and contact-rich behaviors introduce complexities beyond the current framework, presenting an exciting direction for future work.
Appendix A Task Weights
Here, we provide the weights for different costs listed in Sec. 4 in computing classifier guidance for each tasks.
Table 3. Task Weights Table
Task Category Task Task Space Ctrl Static Obs. Dyn. Obs. Way-point Other
Task Space Control Root Path Following 0.2 (root path)
Reaching 0.2 (selected body)
Gamepad Control 0.1 (velocity)
Walk-Perturb 0.1 (forward vel.)
Static Obstacle Avoidance Jump 1 0.06
Forest 1 0.175
Dynamic Obstacle Avoidance Forest + Dynamic 1 1 0.175
Motion In-betweening Motion In-betweening 0.08 Inpainting